How much distilled water will be needed to make an 8 oz. solution that is 80% Alcohol from 100% alcohol?
Answer by ptaylor(2198)

Let x=amt of distilled water needed
Now we know that the amount of pure alcohol that exists before the mixture takes place 1.0(8-x) has to equal the amount of pure alcohol that exists after the mixture takes place(0.8*8). Soooooo
1.0(8-x)=0.8*8
8-x=6.4
-x=-8+6.4
-x=-1.6 oz
x=1.6 oz---amount of distilled water needed
CK
1.0(6.4)=0.8*8
6.4=6.4
Hope this helps---ptaylor